OpenStudy (anonymous):

John's average score on three tests is 87%. What is the least he can score on his fourth test so that his average grade will be at least 90%?

OpenStudy (anonymous):

solve \[\frac{3\times87+x}{4}=90\] fpr x

OpenStudy (anonymous):

\[3\times 87+x=4\times 90\] \[261+x=360\] \[x=360-261=99\]

OpenStudy (anonymous):

hmm doesn't look good for john